OpenStudy (anonymous):

Solve the system below and identify the shape of the graph of the equation. 7y^2+x^2=64 x+y=4

OpenStudy (shamil98):

x = - y + 4 substitute, it's a long process doing this stuff lol. You can tell the first one is a circle from its form and the second equation is a line.

myininaya (myininaya):

The first one is close to a circle but not exactly.

OpenStudy (anonymous):

what do i do with x = -y+4

OpenStudy (shamil98):

Substitute it into the other equation for x. \[\huge 7y^2 + (4-y)^2 = 64\] and solve for y.

OpenStudy (anonymous):

and then find the y (factor y)

OpenStudy (anonymous):

first would I multiply (4-y)(4-y)??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I got 16+y^2, is that correct?

OpenStudy (shamil98):

16 + y^2 - 8y

OpenStudy (anonymous):

oh! I thought one was positive and one was negative!

OpenStudy (shamil98):

(4-y)^2 = (4-y)(4-y)

OpenStudy (anonymous):

so after i get that solution what do i plug it into?

OpenStudy (shamil98):

\[\huge 7y^2 + 16 + y^2 - 8y = 64\] get everything on one side and combine the like terms and then use the quadratic formula or factor it. to get your y values.

OpenStudy (anonymous):

if u confuse, try using this formula : \[\frac{ (-b)\pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

thanks! let me solve first using what shamil gave me :)

OpenStudy (anonymous):

okay I got 1+- 40/2 which gave me 41/2 and -39/2. is that correct? i used the quadratic formula

OpenStudy (anonymous):

did i do this write though if the question says "For questions 10 - 11, solve each system. In your work, identify the shape of the graph of each equation." ??

OpenStudy (shamil98):

7y^2 +16 + y^2 - 8y = 64 8y^2 - 8y - 48 = 0 8(y^2 - y - 6 ) = 0 y^2 -y - 6 = 0 (y-2)(y+3) = 0 y = 2, y = -3

OpenStudy (shamil98):

How did you end up with those values? o_O

OpenStudy (anonymous):

tblue said to use quadratic formula ._. so no?

OpenStudy (anonymous):

it just another way

OpenStudy (shamil98):

(y-3) ( y+2) = 0 oops

OpenStudy (shamil98):

y = 3, y = -2 *

OpenStudy (shamil98):

i like using factoring where possible it's much easier.

OpenStudy (anonymous):

hm, what would i do with those numbers then?

OpenStudy (anonymous):

so u can substitute to x + y = 4

OpenStudy (shamil98):

So anyways. y = 3, y = -2 x+y=4 x + 3 =4 x = 1, x -2 =4 x =6

OpenStudy (shamil98):

So, your solutions are (1,3) and ( 6,-2)

OpenStudy (anonymous):

he's right

OpenStudy (anonymous):

thank you! so @shamil98 (1,3) and (6,-2) would be the equation to each system?

OpenStudy (shamil98):

Those are the solutions to the system.

OpenStudy (anonymous):

& when it says " identify the shape of the graph of each equation" that would be where you said the top eqation was close to being a circle and the bottom equation was a line?